Circular No. 5617

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 2037–404 A. J. Castro-Tirado, S. Brandt, and N. Lund, for the Granat WATCH

collaboration (Danish Space Research Institute, Lyngby; and Space Research Institute, Moscow), report: “On Sept. 23.192 UT, a short-duration x-ray transient was observed by WATCH on Granat. The transient lasted for 110 min and reached a peak intensity of ∼ 1 Crab, though with a somewhat softer spectrum. The position of the transient was α = 20^{h}37^{m}10^{s}, δ = −40^{o}25′.0 (equinox 1950.0). This position is well off the Galactic plane, and quick optical follow-up may be rewarding.”

NOVA SAGITTARII 1992 No. 2 A. D. Scott and A. Evans, Department of Physics, Keele University;

and D. de Martino, International Ultraviolet Explorer (IUE) Observatory, European Space Agency, communicate: “IUE observations of this object have been made on July 31 and Aug. 2. Fine Error Sensor V magnitudes indicate a fading of ∼ 0.08 mag/day since July 20.6 (IAUC 5566): July 31.6 UT, 10.0; Aug. 2.9, 10.2. The ultraviolet spectrum from 120 to 190 nm is characterized by strong emission lines of O I, C I, C II, N I, Si II, Al II, and Fe II, with no evidence of He II (164 nm), C IV (155 nm), or C III] (191 nm). The low-resolution ultraviolet spectrum closely resembles that of the slow nova PW Vul at a similar stage of evolution. Emissions of C II], N II], Fe II, and O III are detected in the low-dispersion spectrum between 190 and 320 nm. P-Cyg profiles have been clearly observed in highdispersion mode, indicating velocities up to 3000 km/s. A reddening of 0.35 ± 0.05 is determined from the 220-nm absorption feature, consistent with the interstellar line equivalent widths observed in high-dispersion mode.”

NOVA CYGNI 1992 C. E. Woodward, University of Wyoming; and R. D. Gehrz, University

of Minnesota, report infrared magnitudes obtained with the 2.3-m Wyoming Infrared Observatory telescope: Sept. 17.24 UT (with bolometer), J = 9.09 ± 0.53, H = 8.83 ± 0.12, K = 8.26 ± 0.09, L = 7.16 ± 0.16, M = 5.85 ± 0.17, N = 4.46 ± 0.29; Sept. 19.10 (with InSb detector), J = 8.44 ± 0.11, H = 8.90 ± 0.05, K = 8.18 ± 0.06, L′ = 6.90 ± 0.06.

1992 September 24 Daniel W. E. Green